# Univ. Lille, Inria, CNRS, Centrale Lille, UMR 9189 CRIStAL, F-59000 Lille, France

### v1

@@ 5 Evaluating CUTECat
We now turn to our experimental evaluation of CUTECat. We first quantify the impact of the optimizations described in Sec. 4, relying on four Catala projects predating this work (Sec. 5.1). We then discuss CUTECat’s ability to identify conflicts (Sec. 5.2), before evaluating CUTECat on the largest Catala program up to this day, namely, an implementation of the French Housing Benefits (Sec. 5.3). All experiments have been performed on a desktop machine featuring an Intel Core i7-12700 running Ubuntu 24.04.1, OCaml 4.14.1 and Z3 4.12.5.
@@ 5.1 Evaluating CUTECat’s Optimizations
To evaluate the performance of CUTECat and its optimizations, we have selected four different, real-world Catala programs as benchmarks for our evaluation. These programs predate our implementation of CUTECat, and have been written by third-party, experienced Catala developers. One benchmark corresponds to the implementation of a fragment of Section 132 of the US Tax Code, defining qualified employee discounts. Other benchmarks implement diverse bodies of law in France: one codifies minimum salary (SMIC), another the computation of a family quotient used in the income tax computation (Family Quotient). The last benchmark computes the housing benefits in the rental case, where additional restrictions have been made: households are located in mainland France and contain less than 10 children. These benchmarks reflect multiple facets of the law, and therefore exhibit diverse encodings in Catala; we provide quantitative details about them in Table 2, namely, the lines of Catala code (including legal specification), the number of default terms and exceptions they contain, and the size of the compiled Python code.
@@
To evaluate the completeness of concolic tools, a standard approach is to measure the code coverage achieved. In our context, this is however hard to perform: the Catala interpreter does not provide any coverage measurements, while results on compiled Catala code, e.g., in Python, are highly unreliable. Indeed, on a small program for which we can manually determine all possible execution paths, a single Python execution reaches a line coverage of 65%, while executing all program paths barely reaches 80%. We suspect that this is due to a bloated translation of default logic into mainstream languages. Our evaluation therefore focuses on analysis times, and on the number of testcases generated. We defer a discussion of the completeness of CUTECat to Sec. 5.2.
| Benchmark | Catala LOC | nb(default terms) | nb(except terms) | Python LOC |
| Section 132 | 133 | 15 | 9 | 343 |
| SMIC | 368 | 17 | 16 | 586 |
| Family Quotient | 776 | 46 | 26 | 1,413 |
| Housing Benefits | 19,655 | 251 | 188 | 31,826 |
| Category | Optims. | Section 132 | SMIC | Family Quotient | Housing Benefits |
| Solver Calls | None | 41 | 138 | 13,224 | 355,848 |
| Solver Calls | All | 24 | 138 | 4,142 | 126,147 |
| Generated Tests | None | 10 | 17 | 381 | 24,995 |
| Generated Tests | All | 10 | 17 | 381 | 17,435 |
@@
We compare in Table 2 the number of solver calls performed and the number of tests generated by CUTECat. These numbers are provided both for naive executions without any optimization enabled and when all optimizations are enabled. The difference in generated tests evaluates the impact of pruning redundant exploration paths through pattern matching case folding; the difference in solver calls represents the additional impact of our optimizations targeting the SMT encoding, i.e., our handling of trivial and unsatisfiable constraints. While these optimizations lead to improvements on almost all benchmarks, except SMIC, their precise impact heavily differs. Avoiding calls to solver when considering trivial or unsatisfiable constraints is almost always beneficial; it reduces the number of calls up to 3.2x for the family quotient benchmark. Conversely, pattern matching folding only helps the housing benefits; other, smaller examples do not rely as much on large pattern matching.
@@
We evaluate the impact of using soft constraints in Table 5; the evaluation is performed with all optimizations enabled. We only consider the Section 132 and housing benefits benchmarks, as our soft constraints target monetary amounts given as inputs, which other examples do not have.
On large, representative Catala examples, we observe that the computational overhead of using soft constraints is small. Additionally, the soft constraints are satisfied for almost all cases, significantly improving the usability of CUTECat. Indeed, 87.04% of the 17,435 generated testcases can be solved with a soft constraint ensuring monetary amounts are multiples of 100, and additionally 9.98% and 2.80% for the multiples of 10 and units, respectively. In total, only 30 of the 17,435 generated testcases (0.17%) do not satisfy any soft constraint. Our encoding of soft constraints particularly benefits from incremental solving: they are only added to a path condition proved satisfiable by the solver, which intuitively corresponds to refining an existing solution. Furthermore, incremental solving significantly reduces solving time, minimizing the impact on CUTECat’s running time of calling the solver several times on a given path. As such, we expect further extensions to soft constraints to have a minor impact on CUTECat’s overall performance.
@@ 5.3 Case Study: Rental Housing Benefits
Our evaluation so far focused on a constrained case of the rental housing benefits, considering only households located in mainland France and with less than 10 children. We chose to enforce these restrictions due to the multiple executions required to evaluate different CUTECat settings, which already required more than one hour each. In this section, we now explore the general, unconstrained case of the rental housing benefits. With all optimizations enabled, CUTECat is able to analyze this program in 6h37m, generating 186,390 testcases. This process generates 1,338,575 solver calls; the total solver time is 366s.
@@
We now compare the overhead of the concolic execution of CUTECat with the interpretation times of all the generated testcases. The reference Catala interpreter evaluates the 186,390 generated tests in 1h29, meaning CUTECat has a 4.5x overhead. While large, this overhead is not unexpected: the concolic interpreter instruments each expression evaluation to collect symbolic constraints, and performs repeated calls to the SMT solver. Prior work on concolic execution observed similar trends: for example, Yun et al. [58, Table 1] measure that KLEE [9] has up to a three order-of-magnitude overhead, while our experimental results are in line with approaches focusing on reducing this overhead [46, 13]. This leads us to believe that CUTECat’s scalability is comparable to concolic tools in other languages; we nevertheless intend to perform a deeper profiling of our interpreter to identify potential bottlenecks.
This comparison also allows us to sanity-check our implementation with respect to the official Catala semantics: we confirmed that CUTECat and the Catala reference interpreter agreed on all generated testcases.
@@ 6 Related Work
Several approaches were previously proposed to reason about default logic. Risch [48] and Cassano et al. [12] both define tableaux-based approaches [23] to reason about propositional formulas expressed in default logic, while Schaub [49] and Linke and Schaub [28] propose query answering methodologies for default logic propositions. By focusing on the underlying logic, these works differ from our approach tailored to program verification, which emphasizes testcase generation and requires reasoning about numerical expressions and program constructs such as structures or enumerations.
In their compiler for the French tax code, Merigoux et al. [36] have relied on coverage-guided fuzzing with AFL [60] to improve the quality of a pre-existing test-suite. Their approach however only generated 275 minimized testcases from a first generation of 30,000 cases. In contrast, CUTECat’s concolic-based approach allows to generate cases exploring different execution paths, while reaching rare ambiguous or unhandled situations.
Monat et al. [39] and de Almeida Borges et al. [1] focus on formalizing date-duration arithmetic and semantics of UTC time respectively, with the former also analyzing date-related ambiguities in legal computations. Our work currently only provides limited support for date operations, but could build upon these formalizations to define suitable symbolic encodings.
@@ 7 Conclusion and Future Work
We are grateful to Rohan Padhye for his suggestions around concolic testing for Catala, and to Patrick Baillot for feedback about earlier versions of this paper. We thank Liane Huttner & Sarah Lawsky for the interesting discussions around the properties this work targets, Louis Gesbert for his technical help around the Catala compiler, and Denis Merigoux for clarifications about the current implementation of default terms in Catala. We appreciated the many discussions and valuable feedback about this work we got from the whole Catala team.

### v2

@@ 5 Evaluating CUTECat
We now turn to our experimental evaluation of CUTECat. We first quantify the impact of the optimizations described in Sec. 4, relying on four Catala projects predating this work (Sec. 5.1). We then discuss CUTECat’s ability to identify conflicts (Sec. 5.2), before evaluating CUTECat on the largest Catala program up to this day, namely, an implementation of the French Housing Benefits (Sec. 5.3). All experiments have been performed on a desktop machine featuring an Intel Core i7-12700 and 128GB of DDR5 RAM (although we have never seen CUTECat use more than 1GB), running Ubuntu 24.04.1, OCaml 4.14.1 and Z3 4.12.5.
@@ 5.1 Evaluating CUTECat’s Optimizations
To evaluate the performance of CUTECat and its optimizations, we have selected four different, real-world Catala programs as benchmarks for our evaluation. These programs predate our implementation of CUTECat, and have been written by third-party, experienced Catala developers. One benchmark corresponds to the implementation of a fragment of Section 132 of the US Tax Code, defining qualified employee discounts. Other benchmarks implement diverse bodies of law in France: one codifies minimum salary (SMIC), another the computation of a family quotient used in the income tax computation (Family Quotient). The last benchmark computes the housing benefits in the rental case, where additional restrictions have been made: households are located in mainland France and contain less than 10 children. These benchmarks reflect multiple facets of the law, and therefore exhibit diverse encodings in Catala; we provide quantitative details about them in Table 2, namely, the lines of Catala code (including legal specification), the number of branching constructs they contain (default terms, if-then-else, pattern matches), and the size of the compiled Python code. We particularly observe that the default terms outnumber the other branching terms, arguing in favor of their importance.
@@
To evaluate the completeness of concolic tools, a standard approach is to measure the code coverage achieved. In our context, this is however hard to perform: the Catala interpreter does not provide any coverage measurements, while results on compiled Catala code, e.g., in Python, are highly unreliable. Indeed, on a small program for which we can manually determine all possible execution paths, a single Python execution reaches a line coverage of 65%, while executing all program paths barely reaches 80%. We suspect that this is due to a bloated translation of default logic into mainstream languages. color=Plum,inline]R: Mention that CUTECat is able to terminate and exhaustively explore all the benchmarks? Our evaluation therefore focuses on analysis times, and on the number of testcases generated. We defer a discussion of the completeness of CUTECat to Sec. 5.2.
Benchmark | Section 132 | SMIC | Family Quotient | Housing Benefits |
Catala LOC | 133 | 368 | 776 | 19655 |
nb(default terms) | 15 | 17 | 46 | 251 |
nb(if-then-else) | 3 | 0 | 24 | 171 |
nb(match) | 2 | 0 | 14 | 139 |
Python LOC | 343 | 586 | 1413 | 31826 |
Category | Section 132 | SMIC | Family Quotient | Housing Benefits | ||||
Optimizations | None | All | None | All | None | All | None | All |
Solver Calls | 41 | 24 | 138 | 138 | 13224 | 4142 | 355848 | 126147 |
Generated Tests | 10 | 10 | 17 | 17 | 381 | 381 | 24995 | 17435 |
@@
We compare in Table 2 the number of solver calls performed and the number of tests generated by CUTECat. These numbers are provided both for naive executions without any optimization enabled and when all optimizations are enabled. The difference in generated tests evaluates the impact of pruning redundant exploration paths through pattern matching case folding; the difference in solver calls represents the additional impact of our optimizations targeting the SMT encoding, i.e., our handling of trivial constraints. While these optimizations lead to improvements on almost all benchmarks (SMIC has no matches or trivial constraints, so it is not affected), their precise impact heavily differs. Avoiding calls to solver when considering trivial constraints is almost always beneficial; it reduces the number of calls up to 3.2x for the family quotient benchmark. Conversely, pattern matching folding only helps the housing benefits; other, smaller examples do not rely as much on large pattern matching. The tables below provide a more detailed analysis of the impact of individual optimizations.
@@
color=Plum,]Some optimizations affect others: trivial is not that interesting in an incremental context? AF: Just by ablation, hard to evaluate. We would need the combination of incr + trivial. I skipped this description
@@
We evaluate the impact of using soft constraints in Table 5; the evaluation is performed with all optimizations enabled. We only consider the Section 132 and housing benefits benchmarks, as our soft constraints target monetary amounts given as inputs, which other examples do not have. On large, representative Catala examples, we observe that the computational overhead of using soft constraints is small. Additionally, the soft constraints are satisfied for almost all cases, significantly improving the usability of CUTECat. Indeed, 87.04% of the 17435 generated testcases can be solved with a soft constraint ensuring monetary amounts are multiples of 100, and additionally 9.98% and 2.80% for the multiples of 10 and units, respectively. In total, only 30/17435 generated testcases (0.17%) do not satisfy any soft constraint. Our encoding of soft constraints particularly benefits from incremental solving: they are only added to a path condition proved satisfiable by the solver – which corresponds to refining an existing solution. Furthermore, incremental solving significantly reduces solving time, minimizing the impact on CUTECat’s running time of calling the solver several times on a given path. As such, we expect further extensions to soft constraints to have a minor impact on CUTECat’s performance.
color=Plum,]add option for script
@@ 5.3 Case Study: Rental Housing Benefits
Our evaluation so far focused on a constrained case of the rental housing benefits, considering only households located in mainland France and with less than 10 children. We chose to enforce these restrictions due to the multiple executions required to evaluate different CUTECat settings, which already required more than one hour each. In this section, we now explore the general, unconstrained case of the rental housing benefits. With all optimizations enabled, CUTECat is able to analyze this program in 6h37m, generating 186390 testcases. This process generates 1338575 solver calls; the total solver time is 366s.
@@
We now compare the overhead of the concolic execution of CUTECat with the interpretation times of all the generated testcases. The reference Catala interpreter evaluates the 186390 generated tests in 1h29, meaning CUTECat has a 4.5x overhead. While large, this overhead is not unexpected: the concolic interpreter instruments each expression evaluation to collect symbolic constraints, and performs repeated calls to the SMT solver. Prior work on concolic execution observed similar trends: for example, Yun et al. [58] measure that KLEE [8] has up to a three order-of-magnitude overhead, while our experimental results are in line with approaches focusing on reducing this overhead [46, 12]. This leads us to believe that CUTECat’s scalability is comparable to concolic tools in other languages; we nevertheless intend to perform a deeper profiling of our interpreter to identify potential bottlenecks. This comparison also allows us to sanity-check our implementation with respect to Catala’s semantics: we confirmed that CUTECat and the Catala reference interpreter agreed on all generated testcases. Furthermore, we can use those same testcases to check whether all Catala backends have the same behavior on the case study.
@@ 6 Related Work
Several approaches were previously proposed to reason about default logic. Risch [48] and Cassano et al. [11] both define tableaux-based approaches [25] to reason about propositional formulas expressed in default logic, while Schaub [49] and Linke and Schaub [30] propose query answering methodologies for default logic propositions. By focusing on the underlying logic, these works differ from our approach tailored to program verification, which emphasizes testcase generation and requires reasoning about numerical expressions and program constructs such as structures or enumerations. In their compiler for the French tax code, Merigoux et al. [38] have relied on coverage-guided fuzzing with AFL [60] to improve the quality of a pre-existing test-suite. Their approach however only generated 275 minimized testcases from a first generation of 30,000 cases. In contrast, CUTECat’s concolic-based approach allows to generate cases exploring different execution paths, while reaching rare ambiguous or unhandled situations. Monat et al. [41] and de Almeida Borges et al. [17] focus on formalizing date-duration arithmetic and semantics of UTC time respectively, with the former also analyzing date-related ambiguities in legal computations. Our work currently only provides limited support for date operations, but could build upon these formalizations to define suitable symbolic encodings. color=Plum,]I had added a small paragraph about LLMs, commented out due to space constraints
@@ 7 Conclusion and Future Work
We thank the anonymous reviewers for their constructive feedback and support of our work. We are grateful to Rohan Padhye for his suggestions around concolic testing for Catala, and to Patrick Baillot for feedback about earlier versions of this paper. We thank Liane Huttner & Sarah Lawsky for the interesting discussions around the properties this work targets, Louis Gesbert for his technical help around the Catala compiler, and Denis Merigoux for clarifications about the current implementation of default terms in Catala. We appreciated the many discussions and valuable feedback about this work we got from the whole Catala team.